# One-Shot Label-Only Membership Inference Attacks

### v1

@@ A.1 Adversarial perturbation magnitude for members and non-members
A.2 Training and implementation details
Training setup.
Implementation details.
The OSLO setup primarily includes the configuration of source and validation models, as well as the parameters for generating adversarial examples. By default, we use Inception [42], ResNeXt50 [44], and PreActResNet18 [46] as the source model architectures. For validation models, we use DenseNet121 [38], VGG13 [43], and ShuffleNetV2 [45] when ResNet18 [13] is the target model, and ResNet18, VGG13, and ShuffleNetV2 when DenseNet is the target model. We follow the assumption that the attacker has no knowledge of the target model architecture and does not include the same architecture in the source/validation models. More combinations of source and validation models are discussed in Appendix A.3. For the adversarial example generation parameters, we set the maximum perturbation to $80/255$ and the number of sub-procedures to 80 (increasing $\epsilon$ by $1/255$ each iteration).
Experimental setup.

### v2

@@ A.3 Effect of combinations of different source and validation models
A.4 Further discussion on threshold $\tau$
A.4.1 Using $\tau$ as the stopping criterion
In OSLO, the generation of adversarial perturbations stops when the confidence of the correct class on the validation model drops below a specified threshold, $\tau$. An intuitive alternative is to stop at the point of label flipping. However, we found that this approach does not guarantee sufficient perturbation to achieve a low FPR. To illustrate this, we compared different stopping criteria, including label flipping, across various thresholds. Table 7 shows the results for ResNet18 on CIFAR-10.
Table 7: Comparison of different stopping criteria on OSLO’s performance on CIFAR-10 using ResNet18.
Stopping criterion TPR FPR
Threshold $\tau=0.1$ 13.1% 3.0%
Threshold $\tau=0.01$ 7.8% 0.9%
Label flip 61.0% 30.8%
As shown in Table 7, a small threshold $\tau$ is required to achieve a low FPR, whereas stopping at label flipping leads to a high FPR. The use of threshold $\tau$ provides better control over the TPR-FPR trade-off, which is crucial for OSLO’s effectiveness.
A.4.2 Choosing $\tau$ in realistic settings
In real-world scenarios, attackers can adapt $\tau$ based on the desired FPR. In OSLO, $\tau$ is calibrated by adjusting the success rate of the transfer-based attack on the target model. For instance, to achieve a 5% FPR, $\tau$ can be set to reach a 95% success rate. This approach may require access to a small set of non-members, which is a common assumption in MIAs [4, 30].
A.5 Effectiveness of OSLO against models trained with adversarial training
Since OSLO relies on adversarial perturbations, increasing model robustness through adversarial training appears to be a natural defense strategy. To explore this, we evaluated OSLO against target models trained with adversarial training.
First, we tested OSLO’s performance when the target models were trained with adversarial training, while the surrogate models were not. The results are presented in Table 8. We observed that adversarial training reduces OSLO’s ability to achieve low FPR, but this comes at the cost of decreased model utility and increased training time.
Table 8: OSLO’s performance against target models trained with adversarial training on CIFAR-10.
Model Threshold $\tau$ Test ACC (%) Training time (min) Attack (%)
TPR FPR
ResNet18 (no adversarial training) 0.01 82.6 23 7.8 0.9
0.001 5.0 0.2
ResNet18 (adversarial training, $\epsilon$=4/255) 0.01 73.8 85 97.8 57.4
0.001 95.6 52.7
We then evaluated OSLO’s performance when both the target model and the surrogate models were trained with adversarial training. The results are shown in Table 9. OSLO remains effective, particularly in maintaining low FPR in this setting. These findings suggest that adversarial training alone may not be an adequate defense against OSLO.
Table 9: Comparison of OSLO’s performance when surrogate models are trained with and without adversarial training.
Target model training Surrogate model training Threshold $\tau$ Attack TPR Attack FPR
Adversarial training ($\epsilon$=4/255) No adversarial training 0.001 95.6% 52.7%
Adversarial training ($\epsilon$=4/255) Adversarial training ($\epsilon$=4/255) 0.001 15.7% 0.7%
A.6 Comparison with YOQO
OSLO shares a similar goal with the recent work YOQO [34] in improving query efficiency for label-only MIAs. Below, we discuss the key differences to highlight our contributions.
Effectiveness at low FPR.
It is widely accepted that MIAs should achieve low FPR to be practical [5, 11, 33]. However, by design, YOQO cannot adjust the TPR-FPR trade-off, making it unable to achieve low FPRs, as noted in their paper [34]. In contrast, OSLO is specifically designed to achieve high TPRs under low FPRs and significantly outperforms existing methods in this metric. Thus, we are not able to have a fair comparison because YOQO cannot achieve the low FPR that OSLO primarily works on.
Methodology.
To the best of our knowledge, OSLO is the first to leverage transfer-based adversarial attacks for label-only MIAs. YOQO, on the other hand, shares a similar approach with LiRA [5], relying on shadow models trained on target samples (IN models) and shadow models not trained on the target samples (OUT models). This approach depends on the similarity between the shadow model and target model architectures and is sensitive to architecture differences [34]. In contrast, OSLO is robust to variations in surrogate model architecture, as demonstrated by our experiments.